OpenStudy (zoeeh127):

What is the solution to the system? 2y= x + 2 3x + 4y=-16

OpenStudy (zoeeh127):

I need someone to explain it to me please do n't just give me an answer

OpenStudy (openstudier):

Do you know how to use substitution or elimination?

OpenStudy (photon336):

for this one it would probably be easier to use substitution

OpenStudy (photon336):

you would need to solve for y in one equation and plug that answer into y in the second equation

OpenStudy (openstudier):

I'm going to use substitution. First, I'm going to simplify the first equation:\[y = \frac{ 1 }{ 2}x + 1\]

OpenStudy (photon336):

=) you're in good hands @zoeeh127

OpenStudy (openstudier):

Then, like @Photon336 said, you would plug that into y for the second equation.

OpenStudy (openstudier):

\[3x + 4(\frac{ 1 }{ 2 }x + 1)= -16\]

OpenStudy (triciaal):

it is much easier not to use fractions. in this case equation 1 you can already isolate x and substitute in eqn 2

OpenStudy (openstudier):

Now, we solve for x:\[3x + 2x + 4 = -16\]\[5x + 4 = -16\]\[5x = -20\]\[x = -4\]

OpenStudy (openstudier):

Next, we take our value for x and plug it into either equation and solve for y.

OpenStudy (openstudier):

\[2y = (-4) + 2\]

OpenStudy (openstudier):

I don't think the fractions are that bad, @triciaal. But, I understand your point.

OpenStudy (triciaal):